OpenStudy (anonymous):

Can someone please help me with this problem. Ive been trying to figure it out ALL DAY and posting it up. 150 freshmen, 125 sophomores, 100 juniors, and 75 seniors must select 4 of its member for student council. How many types of councils are possible if the following condition is satisfied? *I want actual numbers for this problem please! a) One senior is President and one senior is Vice President b) One senior is President and one Junior is Vice President c) The majority of its members are sophomores d) The council has exactly one senior and one freshman.

OpenStudy (anonymous):

what are different posts?

OpenStudy (amistre64):

how many ways to fill the positions right? as opposed to how many ways can the same group of people be selected without regard for position?

OpenStudy (amistre64):

spose there are just 6 choices and 3 positions to fill: 6 pick 3 gives as all the ways to fill the positions: 6*5*4 = 120 possible ways to do it; but, this takes order as seperate counts; ABC is a different count then BCA.

OpenStudy (anonymous):

Yes how many way to fill the positions.

OpenStudy (amistre64):

6 pick 3 ------- is the number of ways to group the people when 3! mixed orders count the same. ABC is the same count as CAB is the same count as BCA etc...

OpenStudy (amistre64):

sen. oth. oth. sen. --- ---- --- --- 75 375 374 74 ; if i see it right for 'a'

OpenStudy (anonymous):

Is this for a and b?

OpenStudy (amistre64):

no; thats just for a :) if i see it right that is; im unsure if we count the seniors in with the others for 2 and 3 or not

OpenStudy (amistre64):

sen jun oth oth --- --- --- --- 75 100 275 274 ; would be my guess for 'b'

OpenStudy (amistre64):

and we multiply these number together

OpenStudy (anonymous):

oh ok. So would B be kinda the same set up?

OpenStudy (anonymous):

?

OpenStudy (amistre64):

similar yes; but I have to read up on if I am actually doing it right, unless you have a way to verify my answers :)

OpenStudy (anonymous):

No I don't have a way to verify these answers

OpenStudy (amistre64):

i think i have an error, the logic being this: total population = all students sen 75 ; total pop is minus 1 student now sen 74 ; total pop is minus 2 student now any total pop -2 any total pop -3

OpenStudy (amistre64):

total pop = 150+125+100+75 = 450 150 fm, 125 sophs, 100 jrs, and 75 srs We want 2 seniors for "a": 75 *74 *448 *447 ............................................ a sr and a jr for "b" 75 *100 *448 *447 ............................................. majority of sophs; at least 3 of them for "c" 125 *124 *123 *447 .......................................... exactly one sr and exactly one fmn for "d" 75 * 150* 225 *224

OpenStudy (anonymous):

I think for a) it is 10 (sen,sen (two given) together with sensen OR ff OR ss OR jj OR fs OR fj OR sj OR sens OR senf OR senj) (order doesn't matter). I don't think the titles are important only the type and you only can select four people. You can do other similar (someone else can).

OpenStudy (anonymous):

?.? Based on the above for a) 75 *74 *448 *447 <== this says the order of the other two members matters.... I agrees with 'estudier' that the order doesn't matter.... so a) 75 *74 *448C2= a huge # -.- b) 75*100*448C2 = another huge # -.-; c) not sure about this one because it's annoying!!! my guess it's = 364831250 how? toooo long to explain @.@ d) 75*150*225C2 = another huge # =.=

OpenStudy (amistre64):

It doesnt say suggest order is important; it just puts the offices in order to fill them. Pres * VP *other *other ; is all it does to make sorting the matter easier to follow

OpenStudy (amistre64):

Pres requires a sen. ; fill it first VP requires a sen. ; fill it next the others are then able to fill in with whatever is left out of th etotal population